### v1

# Quantifying Self-Awareness of Knowledge in Large Language Models

Abstract
Hallucination prediction in large language models (LLMs) is often interpreted as a sign of self-awareness. However, we argue that such performance can arise from question-side shortcuts rather than true model-side introspection. To disentangle these factors, we propose the Approximate Question-side Effect (AQE), which quantifies the contribution of question-awareness. Our analysis across multiple datasets reveals that much of the reported success stems from exploiting superficial patterns in questions. We further introduce SCAO (Semantic Compression by Answering in One word), a method that enhances the use of model-side signals. Experiments show that SCAO achieves strong and consistent performance, particularly in settings with reduced question-side cues, highlighting its effectiveness in fostering genuine self-awareness in LLMs. The code is available online (https://github.com/ybseo-ac/AQE).
1 Introduction
The defining and quantifying of the human-like mental attributes of large language models (LLMs) lie at the heart of a long-standing human curiosity: whether artificial systems can possess minds akin to our own. While the field has demonstrated that LLMs can rival or even surpass human-level performance in rational reasoning (e.g., reading, solving a logical problem) [4, 11, 29], higher-order human-like traits (e.g., self-awareness, personality, emotion) remain poorly understood, despite growing interest [21, 41], partly due to the inherent ambiguity in their definition and measurement.
Among higher-order attributes, self-awareness of knowledge has been considered as a key topic, given its strong association with hallucination detection—a critical factor in ensuring the reliability of LLMs. While hallucination arises from various causes (e.g., failure of reading, reasoning), it is well known that a major cause is when the model is asked questions on the knowledge not pre-trained [37]. Humans possess an innate ability of self-awareness to recognize the possession of their knowledge [13, 17], and thus refuse to answer when they lack the relevant knowledge. In contrast, LLMs lack such self-awareness and therefore tend to produce any probable output even when they lack knowledge, which leads to hallucination. In this respect, self-awareness is a fundamental component in achieving reliable language generation.
Then, how can we define and measure self-awareness of LLMs? Prior work has often equated self-awareness with the behaviour of hallucination detection itself, perhaps because it is the ultimate practical goal we expect from self-awareness. Indeed, recent studies have reported surprisingly high precision on hallucination detection [35, 42, 26, 13, 1], raising hopes that self-awareness in LLMs may be within reach.
However, in this work, we argue that hallucination prediction (i.e., early detection) does not equate to self-awareness. This is because two distinct sources of information are typically involved in the prediction process: (1) information about the model itself and (2) information about the question. As such, hallucination prediction reflects a mixture of model-awareness (self-awareness) and question-awareness. To isolate the true contribution of self-awareness, we propose to disentangle these two components. Through this separation, we offer a rigorous and measurable definition of self-awareness, which we believe serves as a foundational step toward investigating the higher-order mental faculties of LLMs. To enable this separation, we introduce a metric to quantify the question-awareness in hallucination detection, the Approximate Question-side Effect (AQE). The contribution of self-awareness is estimated by subtracting AQE from hallucination detection precision.
Quantifying the contribution of self-awareness to hallucination prediction performance is not merely a matter of curiosity: it carries significant practical implications, in our observation in §3.2. When hallucination prediction relies on question-awareness, models tend to exploit dataset-specific shortcuts to easily achieve high performance. However, such performance fails to generalize when the distribution of questions shifts even slightly. In contrast, relying on self-awareness tends to produce more consistent results across varying question distributions, as it utilizes model-side information rather than question-side cues. To support this claim, in §B, we first analyze dataset-specific shortcut patterns in existing datasets used in previous hallucination detection works. Then we provide the following empirical findings in §3.3 and §5: (1) Datasets containing shortcut cues tend to exhibit higher AQE scores, suggesting stronger question-awareness signals. However, AQE scores drop significantly under shortcut-controlled settings (e.g., out-of-domain scenarios). (2) In such high-AQE datasets, hallucination detection performance shows limited generalizability
Lastly, we also propose a method to enhance the use of model-side information, by leveraging the confidence scores of LLMs more effectively. The proposed method is Semantic Compression by Answering in One word (SCAO). We demonstrate that SCAO performs particularly well in low AQE settings, and the performance is consistent across varying settings of questions.
Our contributions are summarized as follows:
  • We propose a measurable definition of “self-awareness” in LLMs, distinct from question-awareness, and introduce AQE to quantify its contribution. This offers a concrete example for studying human-like mental attributes in language models.
  • We analyze existing hallucination prediction datasets and reveal that much of the reported performance stems from question-side shortcuts, limiting generalizability.
  • We introduce SCAO, a simple one-word answering method that enhances model-side signal utilization, showing strong performance particularly when question-side shortcuts are unavailable and self-awareness becomes essential.
2 Definition
In this section, we first examine the definition of self-awareness of human. Next, we define the task formulation of hallucination prediction, to establish a definition of self-awareness in LLMs. And we review the definitions from previous works. More detailed review on the related works of neuropsychology and LLMs is in §A.
Definition of self-awareness of human In psychology, human self-awareness is defined as the capability of perception of one’s own mental processes or states, which includes thoughts, feelings, emotions, and knowing [28]. In this work, we focus on the self-awareness of knowing certain pieces of knowledge, which is also referred to as self-knowledge [41].
Some studies on LLMs also borrow the term “meta-cognition” from psychology to refer to self-knowledge [20], which may be inaccurate. It is because meta-cognition focuses solely on the conscious level [17], while recent cognitive neuroscience research suggests that the human brain conducts judgment of knowing even at the unconscious level [13], which is even before consciously recognizing the meaning of the question.
Definition of hallucination prediction The term “hallucination” has been broadly used to refer to the phenomenon where a model provides an incorrect answer to a given question [18, 26, 5]. Thus, “hallucination detection” refers to the task of predicting whether a response is incorrect [18, 6].
In this work, we focus specifically on (1) hallucination from factoid questions that examine whether the model possesses certain knowledge, as this has been widely used as a clear and straightforward scenario for exploring hallucination detection [35, 42]. (2) And we focus on an early detection (i.e., prediction) scenario [35, 6, 1] where $k$ is predicted before answer generation, as this setting is more appropriate for examine self-awareness (we describe this in the next section).
To formulate the common problem setting of hallucination prediction, let $\theta$ represent the model, $x$ the query, and $y$ the answer label. The model $\theta$ infers $\hat{y}$ from the input $x$. The correctness of $\hat{y}$ can then be measured by evaluating its similarity to $y$, denoted as $k$, representing a binary value (True/False). Common evaluation methods include string matching [42], GLUE, and G-eval [24]. During this process, a datapoint $s_{\theta,x}=\theta(x)$ is extracted from $\theta$, which contains information about how $\theta$ perceived $x$. We denote this as $s$ for simplicity. Through a series of question-answering and evaluation processes, we obtain a dataset $\mathcal{D}=\{(s_{i},k_{i})\}_{i=1}^{N}$, where $N$ is the dataset size. From this, hallucination prediction is defined as a binary classification task where a learnable module $\phi$ learn to take input $s_{i}$ to predict $k_{i}$, which we note $\hat{k_{i}}=\phi(s_{i})$.
Assuming $\theta$ as a transformer model [38], $s$ can be mainly two forms (1) Hidden state vectors: Transformer models are composed of multiple attention layers, where each layer passes a fixed-size vector (i.e., hidden states) to the next. These vectors encode the semantic and contextual information of the input $x$ [31], and are also known to contain information about the model’s response that will be generated [19], providing a cue for hallucination prediction. (2) Confidence score: This refers to the softmax probability that the causal LLMs predict for the next token generation. While the hidden state is a high-dimensional representation (4096 dimensions in LLaMA-3-8B), the confidence score is a scalar value. As the hidden state contains richer information, it has achieved the best performance as a source for hallucination prediction and has been regarded as the main source [35, 6].
2.1 Formulating the self-awareness of LLM
Figure 1: Pipeline for the prediction of knowing (prediction of hallucination).
We defined hallucination prediction as the process where $\phi$ perceives $s$ to infer $\hat{k}$ to predict $k$. As $s$ represents the state of the model after it has seen a question, it inherently contains information of two distinct objects, the question-side and the model-side, as illustrated in Figure 1. The question-side refers to the objective information that can be shared across different models, such as the domain of question (e.g., science, math, society) and the type of question (e.g., multiple choice, open-ended). For human, this type of information is derived from rational abilities (e.g., classification and reading comprehension), rather than from higher-order mental capacity such as self-awareness [28]. In contrast, the model-side information refers to model-specific attributes, such as the possession of the knowledge that is needed for response, or the degree of confidence for answering. In humans, this corresponds to the domain of self-awareness.
Let us denote the representation of the question-side information in $s$ as $s_{Q}$. If we exclude $s_{Q}$ from $s$, what remains is the information specific to the model $\theta$, denoted as $s_{M}$. Simplifying this, we can rewrite this as $s\approx s_{Q}\oplus s_{M}$, if $s$ is $d$-dim hidden state. By definition, we assume that $s_{Q}$ is the question-side information that is independent of model-side information ($s_{M}$), and thus the two share no overlap. For example, the domain and type of a question remain invariant regardless of the state of the model. Therefore, the previously defined hallucination prediction $\hat{k}=\phi(s)$ can be rewritten as: $\hat{k}=\phi(s_{Q},s_{M})$.
We do not claim that the hidden state $s$ is perfectly linearly separable. However, we adopt this assumption based on the following intuition: Prior studies have empirically shown that the hidden states of transformer models encode multiple properties in a linearly separable manner. For instance, Li et al. [19] demonstrates that the hidden state contains a direction associated with the attribute of “truthfulness”, therefore linearly adding this to the hidden states results in more truthful responses.
When $\phi$ learns to predict $k$ using two sources of information ($s_{Q},s_{M}$), utilizing question-side information can be regarded as question-awareness, while utilizing model-side information can be regarded as model-awareness, which is “self-awareness” in the perspective of model. Thus, self-awareness can be formally expressed as:
$\hat{k}=\phi(s_{M}).$ (1)
[…]
2.2 Definition from previous works
Utilizing $s$ as self-awareness Previous works implicitly regard self-awareness as the act of hallucination detection itself. In other words, the focus has been on the act of predicting $k$ from $s$, with no consideration given to the decomposition of $s$ into $s_{Q}$ and $s_{M}$. As a result, some of the hallucination detection performance reported in those works is partially overestimated by the effect of question-side shortcuts [42, 1]. We analyze such cases in §3.1.
Utilizing $s_{Q}$ as self-awareness In another case, some works define utilization of $s_{Q}$ as self-awareness, which is the opposite concept from our view. [41] defines the term “self-knowledge” as a self-awareness on possession of certain knowledge. And from this definition, they construct the dataset SelfAware to measure the self-awareness capacity. While the definition in the paper is consistent with ours, the construction of the dataset stands on the opposite definition.
The dataset SelfAware consists of “answerable” and “unanswerable” questions, and the capacity of self-awareness is defined as the classification between the two. An unanswerable question refers to one that is philosophical (e.g., "What is a happy life?") or subjective (“Do you like to go to the mountains?”), where no definitive answer can be given, thus inevitably leading to hallucination. This setting is contradicting the term “self-awareness” in three points: (1) the unanswerability defined in the SelfAware involves no model-side information. For example, the question “Do you like to go to the mountains?” is always classified as “unanswerable” in SelfAware, regardless of how much knowledge about mountains is stored in the answerer (e.g., 1B LM, 70B LM, or a human).
(2) The dataset includes fixed labels indicating the unanswerability of each question, therefore unanswerability is entirely independent from the model’s state. As unanswerability represents whether the model can answer a given question, it aligns with $k$ in our problem setting (Equation 1). However, as $k$ in SelfAware is determined entirely by properties of the question and not the model’s state, the task proposed in this dataset becomes $k=\phi(s_{Q})$, excluding $s_{M}$. This problem setting is a question-awareness, not a self-awareness, from our definition. (3) For humans, the ability needed for this classification task is reading comprehension, which is not self-awareness.
3 AQE: assessing question-side effects of hallucination prediction datasets
As question-awareness is using and relying on the question-side information, in this section, we first identify the data-specific shortcuts that cause dependency on the question-side information, through a case study in existing datasets for hallucination prediction. Next, we introduce our novel metric AQE, a method for quantifying the effect of question-side shortcuts in that dataset.
3.1 Case study on question-side shortcuts
We investigate sources of question-side shortcuts in datasets used in hallucination prediction studies. We focus on short-form closed-book factoid question answering scenarios, which include ParaRel [8], Mintaka [34], HaluEval [18], HotpotQA [39], and SimpleQuestion [40]. We describe three main sources of question-side shortcuts, and further in §B.
(1) Broken question The most commonly observed problem across all datasets is insufficient annotations of question and answer pairs. This occurs when questions and answers follow a one-to-many relationship, but the annotations fail to cover them. For example, the ParaRel consists of question-answer pairs such as: “Q: What field does Daniel Bernoulli work in? A: physics”. Although Daniel Bernoulli also worked in the fields of mathematics and medicine (one-to-many relation), only one label is provided, failing to cover all possible correct answers. If an LLM responds with a correct answer but different from the given label (e.g., mathematics), it would still be classified as hallucinated (i.e., $k$ is annotated $False$), which is incorrect. If $\phi$ is trained to predict $k$ using such $\{s,k\}$ pairs, it will likely become biased toward predicting $\hat{k}=False$ whenever the domain of question is “field of work”. This means $\phi$ learns a domain classification task, not self-awareness. While this may improve prediction performance within the dataset, it lacks generalizability. It is because its performance is likely to drop only if the quality of the dataset improves, or if questions are given from unseen domains. As a study of Zhang et al. [42] reports hallucination performance on this dataset, this shortcut might have influenced the reported scores. This issue is also found in other datasets such as SimpleQuestions, which includes instances like "Q: What is a TV action show? A: Genji Tsushin Agedama", that also fail to cover various possible answers.
Figure 2: The portion of $k=True$ for each domain, by LLaMA-3-8B model on the ParaRel train set. The rate is skewed toward 0 or 1 by domain, rather than being centered around 0.5.
The broken question problem is addressed when a question includes detailed constraints that restrict the one-to-many mapping between the question and possible answers. For example, in Mintaka, detailed constraints are added to the questions to ensure a one-to-one mapping (e.g., “Who was the director of The Goodfellas and attended school at New York University’s School of Film?”).
(2) Domain Classifying the domain of a question (e.g., science, society) alone can provide a rough guess of $k$. For example, suppose a model that is extensively trained on science data but is unfamiliar with society or history domains. Since the model is likely to make more hallucination when given questions in the history domain. In this way, $k$ can be biased toward $True$ or $False$ by domains, as shown in Figure 2. In such cases, the task of predicting $k$ with $\phi$ is more a domain classification, which is question-aware and not model-aware.
Table 1: The portion of correct answer for each question type, by LLaMA-3-8B. The rate for binary type is in bold.
(a) HotpotQA
Type $p(k=True)$
Bridge 0.6828
Comparison 0.8477
(b) Mintaka
Type $p(k=True)$
Entity 0.5508
numerical 0.4011
date 0.5968
string 0.6315
boolean 0.7283
When we consider a case of a human, it becomes clear that such shortcuts have limited effectiveness and are far from self-aware. For example, suppose this model is not familiar with the history domain but still has knowledge about Abraham Lincoln. If this model were human, he could easily “feel” his inner state and recognize that he possesses knowledge about Lincoln, despite not being familiar with other historical issues. However, if the model were an LLM and $\phi$ is trained to exploit only question-side shortcuts, $\phi$ would tell the model does not know about Lincoln, though the model actually possesses the knowledge of Lincoln. Therefore, while the domain information of a question can provide a naive approximation of $k$, relying on it imposes limitations on precision. This again highlights that utilizing self-awareness is the ultimate direction in precise hallucination prediction.
(3) Question type Question type (e.g., short-answer, multiple-choice) also provides a strong hint for predicting $k$. The average probability of $k=True$ (denoted as $p(k=True)$) for binary-choice questions is at least 0.5 for random choice, significantly higher than for open-ended questions with 0 for random choice, as described in 1(b). In such a case, $\phi$ may learn a shortcut where it automatically predicts $k=True$ whenever it detects a binary-choice question. HotpotQA, HaluEval, and Mintaka contain such binary-choice questions.
There are various other question-side shortcuts, which are described in the §B. These shortcuts can be identified by considering various scenarios in which they may act as shortcuts, and it is likely that some shortcuts remain undiscovered, as it is very subtle to determine. Therefore, manually identifying and removing them from datasets is nontrivial. That is why we introduce AQE in the next section, a method for approximately assessing the total effect of question-side shortcuts without human investigation.
3.2 Approximate question-side effect
In this section, we describe the concept of AQE. Repeating §2, the model $\theta$ is given a question $x$ to generate answer, where the correctness is denoted as a binary representation $k$. A hidden state $s$ is extracted at the first position of the answer. And $s$ is consist of model-side $s_{M}$ and question-side information $s_{Q}$. Learnable module $\phi$ can perceive $s$ to infer $\hat{k}$ to predict $k$, which is hallucination prediction. Self-awareness is defined as utilizing $s_{M}$ to predict $k$. Here, what we ultimately aim to measure is the effect of utilizing $s_{M}$ in predicting $k$, denoted as $\mathcal{A}(\phi(s_{M}))$. $\mathcal{A}(\cdot)$ is a metric that measures the correctness of the predicted $\hat{k}$. As all we can measure is $\mathcal{A}(\phi(s_{Q},s_{M}))$, we decompose this as follows: $\mathcal{A}(\phi(s_{Q},s_{M}))\approx\mathcal{A}(\phi(s_{Q}))+\mathcal{A}(\phi(s_{M}))$ This allows us to measure the effect of self-awareness as follows:
$\mathcal{A}(\phi(s_{M}))\approx\mathcal{A}(\phi(s_{Q},s_{M}))-\mathcal{A}(\phi(s_{Q}))$ (2)
AQE as a Shapley analysis
To provide a clearer theoretical ground, we formalize AQE as a special case of Shapley analysis [fryer2021shapley], which evaluates the impact of individual factors on the outcome. Specifically, AQE corresponds to the concept of marginal contribution—a metric that quantifies the separate contriubtion of a single factor.
$\Gamma_{\beta}(\alpha)=\Gamma(\alpha\cup\beta)-\Gamma(\beta)$ (3)
Equation 3 is the general formulation of marginal contribution. Here, $\alpha$ and $\beta$ represent individual components of a system. $\alpha\cup\beta$ represents a state in which these components are mixed together. $\Gamma(\cdot)$ represents the baseline performance metric (e.g., AUROC). $\Gamma_{\beta}(\alpha)$ quantifies the impact of removing $\beta$, thus isolating the contribution of $\alpha$. In our setting, $\alpha$ and $\beta$ correspond to the model-side and question-side information, respectively. $\alpha\cup\beta$ can be interpreted as the information contained in the hidden state of the model, which integrates both the model and question side information. $\Gamma(\alpha\cup\beta)$, $\Gamma(\beta)$, and $\Gamma_{\beta}(\alpha)$ correspond to $A(\phi(s))$, $A(\phi(s_{Q}))$, and $A(\phi(s_{M}))$, respectively.
Computing of self-awareness
Computing $A(\phi(s_{Q}))$ is achieved by introducing a distinct model $\theta^{\prime}$ (where $\theta^{\prime}\neq\theta$) which is optimized to only embed basic properties of the input question (e.g., domain or question type), $s^{\prime}_{Q}\approx s^{\prime}=\theta^{\prime}(x)$. A representative example of $\theta^{\prime}$ is sBERT [31]. sBERT is a very small model with only 22.7M parameters, but it is optimized to generate an embedding vector $s^{\prime}$ from input text $x$ (e.g., question). sBERT is known to capture high-level information from text as effectively as $\theta$ with a larger architecture (e.g., LLaMA-3-8B), as long as the target information is simple enough (e.g., domain classification). Therefore, while $s_{Q}^{\prime}$ and $s_{Q}$ reside in different representational spaces of two distinct models, they are assumed to capture similar high-level information ($s_{Q}^{\prime}\sim s_{Q}$). Conversely, due to its small size, we can assume $\theta^{\prime}$ holds a very small amount of knowledge, which makes the knowledge distribution of $\theta^{\prime}$ and $\theta$ independent ($s^{\prime}_{M}$ and $s_{M}$ are independent). This makes $s^{\prime}_{M}$ ignored when $\phi$ learning to predict $k$ (correctness of $\theta$) from $s^{\prime}$ (hidden state from $\theta^{\prime}$).
This results in following equation, where $\phi$ and $\phi^{\prime}$ share architecture but are trained independently, as $s_{Q}$ and $s_{Q}^{\prime}$ lie in different representational spaces.
$\mathcal{A}(\phi(s_{Q}))\approx\mathcal{A}(\phi^{\prime}(s^{\prime}))$ (4)
The resulting $\mathcal{A}(\phi(s_{Q}))$ of Equation 4 is defined as AQE. To summarize, $\theta^{\prime}$ predicts $k$ of $\theta$ without using information of $\theta$. Intuitively, a distinct model $\theta^{\prime}$ predicts whether $\theta$ will succeed on a given question, using only the information from the question. As no information about $\theta$ is involved, this setup excludes self-awareness and reflects only question-awareness. Together with Equation 2, we can derive $\mathcal{A}(\phi(s_{M}))$, the component of self-awareness in the measured hallucination prediction performance: $\mathcal{A}(\phi(s_{M}))\approx\mathcal{A}(\phi(s))-\mathcal{A}(\phi^{\prime}(s^{\prime}))$
However, this formulation of AQE holds only under the assumption that $s$ is the hidden state format. When $s$ is a confidence score, AQE cannot be directly applied because confidence score is a one-dimensional value, which is too saturated to embed high-level information of the question.
Table 2: AQE assessment on datasets. Prediction of $k$ from LLaMA-3-8B, with $s^{\prime}$ from sBERT.
short-form long-form
ParaRel Mintaka HaluEval HotpotQA SimpleQuestion Explain
$p(k=True)$ 54.31 55.01 37.51 32.71 19.08 31.63
$p(k=False)$ 45.68 44.98 62.48 67.28 80.19 68.36
AQEacc 73.26 63.50 68.55 72.50 82.36 65.65
AQEauc 82.61 66.67 68.37 70.14 68.13 69.40
3.3 Measuring AQE across datasets
In this section, we measure AQE across hallucination prediction datasets. The model $\theta$ is LLaMA-3-8B-Instruct [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B, and we evaluate it mainly on short-form factoid datasets (e.g., ParaRel, Mintaka, Halueval, HotpotQA), and additionally on long-form factoid datasets (e.g., Explain). The details of each dataset are provided in the §D. We use two metrics: AUROC and accuracy. AQE for each metric is denoted as AQEauc and AQEacc, respectively. We also report the $p(k=True)$ and $p(k=False)$, the bias of binary label $k$. We present results for a larger (70B) model in §E, which shows similar trends.
As shown in Table 2, the AQEauc for most datasets is close to or exceeds 0.70, rather than centering around 0.5. This indicates that many datasets exhibit a strong effect of question-side shortcuts, and a model can easily achieve AUROC over 0.70 without any self-awareness (i.e., perception of model-side information), relying solely on question-aware skills such as domain classification. Though previous works [35, 42] have reported to achieve AUROC over 0.80 in hallucination prediction using these datasets, it could be a statistical artifact that is hard to generalize. In all datasets, AQEacc is consistently higher than the $p(k=True)$, indicating that AQE captures question-side effect beyond just the bias of the binary label.
3.4 AQE in refined dataset
In §3.1, we analyze that the type and the domain of a question can act as question-side shortcuts in predicting $k$. Fortunately, some datasets (HotpotQA, Mintaka, and ParaRel) include tags for this information, which allows us to control for it. We analyze AQE for each dataset before and after this control. In Table 9, the “+type” column refers to the dataset after excluding binary-type questions. “+domain” indicates a regrouping of the train and test data according to their categories, such that the domains do not overlap (i.e., out-of-domain setting). The refinement process is detailed in the §D.
Table 3: AQE score of dataset and LLaMA-3-8B model. The version (original, type , domain) with the lowest AQE within each dataset is highlighted in bold.
Mintaka HotpotQA ParaRel Explain
original +type type + domain original +type original +domain original +domain
$p(k=True)$ 55.01 49.71 53.07 42.33 29.12 54.31 60.45 31.63 39.83
$p(k=False)$ 44.98 50.28 46.92 62.48 70.87 45.68 39.54 57.66 60.16
AQEacc 63.50 59.81 59.04 68.55 76.03 73.26 55.09 65.65 61.21
AQEauc 66.67 64.06 61.62 68.37 55.37 82.61 57.55 69.40 61.89
The experimental results can be summarized in two main points. (1) Applying refinement leads to a significant reduction in AQE. This demonstrates that through post-processing, we can get a dataset with lower dependency on the question-side information, which is more suitable for evaluating self-awareness. (2) AQE still remains even after refinement. This suggests the presence of a question-side effect that we have not yet identified or controlled for.
4 SCAO: an approach to enhance the use of $s_{M}$
In this section, we propose Semantic Compression through Answering in One word (SCAO), a hallucination prediction method designed to maximize the utilization of $s_{M}$. While the $\phi$ predicts $k$ from input $s$, which consists of both $s_{M}$ and $s_{Q}$, $\phi$ tends to depend more on $s_{Q}$ which offers an easier shortcut, as discussed in §3.1. To prevent this problem, we need a method that strengthens the preference of $\phi$ for $s_{M}$.
To address this, we focus on the confidence score as a source of $s$. In §3.2 we noted that the confidence score is a 1-dimensional scalar where information is extremely saturated, which in turn makes it unlikely to carry high-level information of the question ($s_{Q}$). This makes the confidence score closer to $s_{M}$ ($s\approx s_{M}$). Therefore, using the confidence score alone or aggregating it with the hidden state may increase the model’s dependency on $s_{M}$.
Figure 3: Structural analogy between 1) dense retriever and 2) causal LM.
However, as the confidence score is highly saturated and carries limited information, an approach is needed to amplify the information contained in it. SCAO is a method designed to more effectively express confidence in knowledge that can be represented as entities. The approach is straightforward: we insert a system instruction before the question, prompting the model to “you must answer in one word”. The rationale behind why this technique can improve the use of $s_{M}$ is as follows.
Causal LLM is an entity retriever when compressed The confidence score is calculated by the maximum inner product score between the last hidden state of $\theta$ and the token embedding vectors (i.e., decoder head). We analyze that this structure is analogous to the maximum inner product search (MIPS) used in dense retrieval [16] (Figure 3). And we focus on the calibration function of similarity scores in dense retrieval. Previous research on dense retriever systems such as Faiss [7] leverages this calibration through a range search, which finds all the document vectors that are within some distance threshold.
This concept can be interpreted in reverse that we can evaluate whether knowledge is within the vector DB, with a fixed confidence threshold. For example, suppose that a vector database contains documents about the biography of Newton but contains rare data about the biography of Lincoln. Querying “Give me an explanation on Lincoln” may result in few documents with confidence scores over the threshold. In contrast, querying "Give me an explanation on Newton" would likely retrieve a greater number of documents exceeding the threshold, reflecting a stronger alignment between the query and the knowledge contained in the database.
Figure 4: Y-axis is the top-7 candidates of the first token of the answer to the question “Please give me an explanation about Breaking Dawn”. The X-axis is the probability for each candidate. Left is for one-word prompt, and the Right is for normal prompt.
However, a causal LM not only performs entity retrieval but also generates full sentences by connecting these words. Such considerations of grammatical words and sentence structure may act as noise for a calibration signal. Therefore, by minimizing the consideration of sentence structure in the model, the LLM will become more analogous to an entity retriever. If LLM becomes more analogous to an entity retriever, its behavior will become more similar to the calibration properties in dense retrieval. A straightforward way to minimize the consideration of grammatical context in the model is to instruct it to “you must answer in one word”, under the assumption that the model is well-trained to follow instructions.
In Figure 4, we show the confidence pattern at the first token of the answer, with and without SCAO when the model is provided a question “Please give me an explanation about Breaking Dawn”. With the one-word prompt (Left), the model appears to attempt to retrieve knowledge related to "Twilight," which is the series name of Breaking Dawn. In contrast, with the normal prompt (Right), the model tends to repeat the question entity, "Breaking Down". Since it chooses the easier path, the overall probabilities are higher. Further analysis of the confidence pattern is in §C.2. In §5, we empirically show that applying SCAO leads to improved hallucination prediction performance, especially more in settings with lower AQE dataset, where the use of model-side information becomes more critical.
5 Experiment on hallucination prediction approaches
In this section, we evaluate hallucination prediction approaches across multiple datasets and their refined versions.
Approaches Previous approaches for hallucination prediction can be broadly categorized into three. (1) confidence-based: this utilizes the softmax probability of the answer token [9]. It is utilized in other forms, such as perplexity [32] or energy [23]. We adopt a simplified method that takes the mean of top-$n$ softmax probabilities of the first answer token and applying a threshold. $n$ and the threshold $t$ are learnable $\phi$. (2) hidden-state-based: This approach feeds the hidden-state vectors of a model $\theta$ into learnable $\phi$. $\phi$ can be a linear layer [19] or more complex architecture [1, 6]. We adopt a three-layer deep neural network. (3) aggregation: This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ [35]. In Table 10, Conf, Probe, and Conf + Probe represent the three evaluated approaches, respectively. Detailed explanations are provided in §E.1. We conduct experiments using instruction-tuned LLaMA models of two different sizes (8B, 70B).
Dataset and metric For the evaluation of hallucination prediction, we narrow down our focus to the datasets from §3.3 that support refinement (Mintaka, HotpotQA, ParaRel, Explain). We also report the gap between the metric performance and AQE, denoted as $\mathcal{A}(\phi(s_{M}))$. However, we do not report the $\mathcal{A}(\phi(s_{M}))$ for the methods that use only the confidence score (e.g., Conf), because they lack question-side information, making it impossible to compute $s_{Q}$. In Table 10, original refers to the unrefined version of dataset, while + type and + domain indicate versions refined based on question type and domain, respectively. Detailed descriptions are in §D.
Results We observe the following points. (1) On the datasets of original, the performance shows very promising records of around AUROC 0.80, the highest among all versions. However, the $\mathcal{A}(\phi(s_{M}))$ is the smallest, suggesting that the performance is largely attributed to question-side shortcuts. (2) In refined versions (+ type, + domain), the performance measures sharply drop. For example, the AUROC for HotpotQA drops from 80.58 to 73.17 after refinement. This again demonstrates that the performance reported in previous works was largely driven by shortcuts.
Table 4: Hallucination prediction performance (AUROC) of instruction-tuned 8B and 70B LLaMA models across multiple datasets.
(a) Mintaka
8B 70B
original + type + type + domain original + type + type + domain
auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$
Conf 64.88 - 64.61 - 69.23 - 69.41 - 67.16 - 66.35 -
Conf (SCAO) 72.13 - 72.01 - 75.51 - 73.64 - 72.78 - 72.46 -
Probednn 77.09 8.54 75.70 11.64 72.79 11.17 76.43 10.68 75.54 10.19 71.03 12.26
Conf + Probe 79.10 10.55 77.54 13.48 70.77 9.15 78.93 13.18 77.68 13.33 70.80 12.03
Conf + Probe (SCAO) 79.41 10.86 77.89 13.83 74.89 13.27 78.86 13.11 77.22 12.87 72.89 14.12
(b) HotpotQA
8B 70B
original + type original + type
auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$
Conf 74.88 - 68.82 - 73.33 - 72.87 -
Conf (SCAO) 77.70 - 73.81 - 74.13 - 73.42 -
Probednn 80.58 12.21 73.17 17.80 74.41 10.63 69.42 14.86
Conf + Probe 81.08 12.71 73.87 18.50 77.33 13.55 73.06 18.50
Conf + Probe (SCAO) 83.39 15.02 75.51 20.14 77.28 13.50 73.52 18.96
(c) ParaRel
8B 70B
original + domain original + domain
auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$
Conf 71.03 - 59.51 - 70.87 - 59.92
Conf (SCAO) 69.23 - 73.12 - 73.45 - 74.51
Probednn 88.76 6.15 70.34 12.79 89.88 2.90 68.66 15.37
Conf + Probe 88.78 6.17 73.08 15.53 90.01 3.03 68.86 15.57
Conf + Probe (SCAO) 88.95 6.34 76.09 18.54 89.82 2.84 70.84 17.55
(d) Explain
8B 70B
original + domain original + domain
auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$
Conf 49.45 - 50.57 - 48.96 - 46.81 -
Conf (SCAO) 62.90 - 62.28 - 57.21 - 59.54 -
Probednn 84.68 15.28 68.55 6.66 83.69 16.00 66.58 8.87
Conf + Probe 84.89 16.49 69.15 7.26 83.68 15.99 66.78 9.07
Conf + Probe (SCAO) 85.42 17.02 70.04 8.15 84.94 17.25 68.67 10.96
(3) However, the AQE gap is larger in the refined datasets, indicating that when question-side effects are reduced, the use of model-side information becomes more activated. (4) Methods that rely solely on the confidence score (Conf(SCAO)) perform poorly on the original datasets, but show smaller performance variation across different data versions. And in some refined settings, it even outperforms other baselines. It is somewhat counter-intuitive that Conf(SCAO) outperforms hidden-state based methods, though it is provided significantly smaller amount of information. This suggests that the confidence score with SCAO captures a substantial amount of $s_{M}$, contributing to its strong generalization performance. (5) The Conf + Probe (SCAO) shows the largest $\mathcal{A}(\phi(s_{M}))$ across all refined versions of the datasets, suggesting a stable and effective direction for achieving self-awareness. (6) These trends remain consistent across models of different sizes and also hold under the accuracy metric §E.4.
(7) Even in a long-form question answering scenario (Explain), similar trends are observed, including the superiority of SCAO. This suggests that the “answer in one word” instruction remains effective even in scenarios where a one-word answer is inherently infeasible.
6 Conclusion and limitation
In this work, we provide a measurable definition of self-awareness in LLMs, distinguishing it from mere hallucination prediction by isolating model-side and question-side effects. We introduce AQE to quantify the question-side cues, enabling a clearer attribution of self-awareness. This definition and measurement offer a reference point for studying human-like mental traits in LLMs. Empirically, we show that much of the prior performance in hallucination prediction is driven by question-side shortcuts rather than self-awareness, and that our proposed method, SCAO, enhances self-awareness and yields more robust and generalizable predictions across diverse settings.
However, applying our framework beyond knowledge-recall tasks to more complex settings (e.g., reasoning, text reading, and long-form open-ended questions) remains an important direction for future work.
Appendix A Related works
As human self-awareness has been extensively studied in cognitive psychology and neuroscience, its core mechanisms can be leveraged to structure and categorize approaches for evaluating internal confidence and hallucination in LLMs.
Self-awareness in humans: Insights from cognitive neuropsychology
Extensive research in cognitive psychology and neuroscience has shown that human self-awareness—particularly in the context of knowing whether one knows something—relies on layered cognitive processes. According to studies such as [17, 13, 3], two key mechanisms underpin this self-assessment.
1) Unconscious level: When a query is received, the brain initiates implicit memory retrieval, evaluating whether candidate memories fit the contextual cues. This process activates regions such as the orbitofrontal and prefrontal cortices within 300–500 milliseconds [33, 13], distinct from areas responsible for linguistic output, such as the posterior temporal lobe (i.e., Wernicke’s area) [2]. Several factors—such as the amount, accessibility, and vividness of retrieved information—act as internal cues signaling potential knowledge [17].
2) Conscious level: The results of these unconscious processes are then consciously evaluated through metacognitive strategies. These include checking for logical and temporal consistency or aligning retrieved information with known frameworks. This layered evaluation reflects human self-awareness—our capacity to introspect on our own knowledge states and confidence levels.
The dual-process theory [15] offers a broader framing: rapid, intuitive processes dominate simple recall tasks, while complex, deliberative processes support tasks such as reasoning or problem-solving.
Self-awareness in LLMs: A perspective on hallucination detection
In large language models (LLMs), the concept of self-awareness can be reinterpreted as the model’s ability to internally assess whether it possesses sufficient knowledge to answer a question accurately. This introspective capacity—whether performed before or after answer generation—is directly related to hallucination detection mechanisms. Our analysis proposes aligning hallucination mitigation strategies with the structure of human self-awareness processes.
1) Before-generation: Approaches that attempt to detect potential hallucinations *before* the model generates an answer mirror the unconscious processes of humans. These methods, including our own, rely on internal indicators such as activation patterns or uncertainty proxies to determine knowledge sufficiency prior to verbalization [25]. Benchmarks like Mintaka [34] and ParaRel [8] primarily test immediate factual recall, aligning well with this layer of self-assessment.
2) After-generation: Other approaches evaluate the model’s response after generation, resembling conscious-level reasoning in humans. These include multi-pass generation, self-consistency checks, and the integration of external tools such as retrievers [5, 26, 6]. Benchmarks requiring structured reasoning, such as MMLU [12], TruthfulQA [22], and ELI5 [10], are more aligned with this stage, as they demand deliberative thought processes rather than pure retrieval.
Importantly, not all hallucinations can be attributed to failures in self-awareness. For instance, hallucinations in open-book QA tasks may arise from comprehension or inference errors rather than epistemic uncertainty. Benchmarks like SQuAD [30] and FEVER [36] exemplify this issue, being more about information grounding than self-assessment.
Overall, hallucination in LLMs encapsulates a variety of cognitive failures. Addressing them requires different forms of internal awareness—ranging from assessing memory sufficiency to verifying logical coherence. Ultimately, a robust system may need to combine multiple self-assessment mechanisms. In this work, we focus on introspective strategies relevant to knowledge recall before answer generation, drawing inspiration from unconscious-level self-awareness in humans.
Appendix B Other question-side shortcuts
Answerability
Some questions are inherently unjudgeable in terms of correct or incorrect. This includes preference-based questions, hypothetical scenarios, and philosophical inquiries. For such questions, any answer could be considered correct or hallucinated, depending on the perspective of the evaluator. For example, if an LLM is asked “What color do you like?” and responds “Blue,” the correctness of this answer depends on the interpretation: 1) If correctness is judged based on contextual appropriateness, the answer is valid. 2) If correctness is judged based on the idea that LLMs cannot have personal preferences, the answer could be labeled as hallucinated. In this case, $\phi$ can exploit this pattern by simply identifying unanswerable questions and assigning a fixed label (either $k=0$ or $k=1$) across all such cases. This makes the hallucination detection process dependent on question-awareness, rather than assessing self-awareness.
The dataset SelfAware collects only unanswerable questions and categorizes them into different types (e.g., no scientific consensus, imagination, completely subjective, too many variables, philosophical). However, distinguishing these does not require self-awareness. Instead, it is primarily an act of reading comprehension, where the model identifies the nature of the question rather than assessing its own knowledge state.
Time-sensitive question
Time-sensitive questions are inherently difficult for LLMs to answer accurately, as LLMs lack a robust understanding of time [14]. As a result, questions involving temporal information will be biased toward hallucinated responses. Datasets such as HaluEval, Mintaka, and TruthfulQA include such questions (e.g., "How old is Barack Obama?", "When did the most recent pandemic occur?").
Complexity
Complexity awareness is another question-dependent approach to estimating $k$. This aligns with the case of “too many variables” in answerability awareness, where, if a question is too difficult, the model is more likely to fail, making it advantageous to predict "unknown" by default. However, the notion of complexity is relative. If a model has extensive knowledge in a certain domain, it may still answer correctly even if the complexity is high. These attributes may connect complexity with category awareness. Additionally, our analysis suggests that questions within a single dataset tend to have similar levels of complexity. Therefore, distinguishing $k$ based on complexity within a dataset is expected to be relatively rare in the general experimental setup.
Appendix C Details on SCAO
C.1 Efficiency of first token as a discriminator
Figure 5: Probability pattern of the hallucinated answer, by LLaMA3-8B. Each bar stands for the probability (0,1) of the corresponding token.
Previous works on confidence-based hallucination detection research mostly utilize the confidence score of all tokens in the answer sentences, with normalization such as averaging [6]. Utilizing more information is ultimately more advantageous; however, it also has several drawbacks. We observe a pattern that as the entity name length increases, the average confidence tends to rise. For example, Figure 5 depicts the confidence pattern of the hallucinated question-answer pair “Question: Give me an explanation about Obama. Answer: Harry Potter and the Philosopher’s Stone”.
Up to the token “Harry Potter”, the confidence is near zero since it conflicts with the question. However, from a “philosopher”, confidence increases to a near maximum, as the previous context of “Harry Potter” supports it strongly. Thus, the average confidence tends to increase regardless of whether it makes sense, when the entity name gets longer or the sentence contains more grammatical elements. This observation is supported by the analysis in Figure 6 (Left), which shows that the correlation between the mean confidence and the $k$ tends to decrease as the token increases.
Figure 6: Y-axis is a correlation between the mean confidence and the $k$. The X-axis of each figure stands for (Left) the number of tokens from the start point of the answer, and (Right) the number of candidates used to calculate the mean. The LLaMA3-8B and (s,k){(s,k)} datasets from Mintaka are utilized.
We also observe that averaging the confidence scores across top-$n$ vocabulary candidates, rather than just the top-1, shows a stronger correlation with the label $k$, particularly peaking around nn=15 (Figure 6 (Right)). This suggests that incorporating more samples of distance provides more information about the relationship between the output vector and the token space.
C.2 Confidence pattern of SCAO
Table 5: The number and portion of each case, when questions from the test set (total 2152) of Explain are asked to the LLaMA-3-8B-Instruct model using various prompts. The columns represent each prompt style. In the rows, “repeating subject” refers to cases where the top-1 candidate for the first token of the answer is a component of the queried subject entity. “The” refers to cases where the top-1 token is "the."
one-word prompt normal prompt
repeating subject 1819 (84.5%) 261 (12.1%)
"the" 383 (17.7%) 5 (0.2%)
In Figure 4, we show how the model reacts at the first token of the answer in both one-word prompts and the normal prompt.
First, in non-compressed cases (queried with a normal prompt), the following patterns are frequently observed: (1) The response often starts by repeating the entity name mentioned in the query. (2) The response begins with grammatical function words such as "The" or "A". In other words, the model tends to take the easy path. As a result, the probability of the initial token is generally inflated, regardless of whether the model truly knows the subject.
On the other hand, when prompted to answer with a one-word response, the first token often corresponds to the initial token of a word encapsulating the entity’s characteristics. For example, in response to the question "Please give me an explanation about ’Breaking Dawn’.", the first candidate token was "Tw" (the first token of "Twilight"). In other words, with one-word prompting, the model shows a stronger tendency to retrieve its own knowledge related to the entity. This trend is also reflected statistically. Among the 2152 test samples in the Explain dataset, the case where the top-1 candidate of the first token of the response is a component of the entity is 84.5% for normal prompting, significantly outpacing the 12.1% for one-word prompting. Similarly, the first token being "the" occurred in 17.8% of normal prompting cases, compared to just 0.02% for one-word prompting. (Table 5)
C.3 Rationale on why confidence-based method is better in generalization
In §5, we observed that the confidence-based method (SCAO) outperforms the hidden-state-based method (probing) in the out-of-domain setting. This result is counter-intuitive, as confidence scores are highly saturated scalar values, whereas hidden states are high-dimensional vectors capable of carrying much richer information. We suggest the following rationale for this result, examining how the probing and SCAO learn to predict hallucinations
SCAO and probing are fundamentally similar. Probing directly utilizes the raw hidden state of $\theta$, while SCAO focuses on the last hidden state of $\theta$, which is projected onto the vocab embedding space.
Let us assume a knowledge space (denoted as $S_{k}$) (Figure 7), which represents the embedding of each knowledge in the $\theta$. And we term the gray area in the $S_{k}$ as a boundary of knowing of $\theta$, which represents the area where $k=True$ (model possesses the knowledge). This space is hypothetical and unknown but needs to be discovered to predict hallucination of $\theta$. To approximate this, what we have at hand are (1) the 4096-dimensional (in the case of LLaMA-3-8B-Instruct) hidden states (denoted as $S_{h}$) and (2) a vocab embedding space (denoted as $S_{e}$) of the same dimension, with vocab embedding vectors (denoted as $v$) distributed across $S_{e}$.
In probing, a linear layer is trained to map $S_{h}$ to $S_{k}$. The weight of the linear layer is supposed to be a direction vector that represents a principal component of the boundary of knowing. Thus, an inner product with this vector tells if the given hidden states match the direction. Since it utilizes all 4096 dimensions to describe $S_{k}$, it offers high informational resolution, leading to generally strong performance.
Figure 7: Illustration on two methods (probe, SCAO) approximating the boundary of knowing of $\theta$. In $S_{e}$ (lower right), the green balls are the last hidden state vector that is mapped to the vocab space. SCAO learns the threshold of distance between the hidden state and $v$ to classify $y$ of each ball.
Conversely, SCAO assumes that $S_{e}$ approximately aligns with $S_{k}$, which means the $v$ (gray balls in Figure 7) aligns with the boundary of knowing (gray area in Figure 7). SCAO figures the shape of $S_{k}$ by measuring the distance between the hidden state vector from the last layer (green balls in Figure 7) and embedding vectors $v$. These mechanisms yield the following properties: (1) SCAO leverages $S_{e}$ and $S_{h}$, thus utilizing more information than probing, which only uses $S_{h}$. (2) However, this information is compressed into a single scalar value, distance, leading to lower information resolution, showing lower performance than the probe. 3) Despite the lower resolution, this simplification appears to enhance generalization. For instance, in out-of-domain scenarios, probing struggles with unfamiliar features in $S_{h}$, while SCAO effectively handles these novel features by employing its simplified distance-based measure.
Since probing and SCAO reflect slightly different aspects of $S_{k}$, combining these two methods in a feature fusion appears to provide an additional performance boost by leveraging their complementary strengths.
Appendix D Detail of datasets
D.1 Datasets and their refinement strategies
In this paragraph, we present details about the benchmark dataset for evaluating the hallucination prediction method: Mintaka [34], ParaRel [8], HaluEval [18], HotpotQA [39], and Explain. We also describe the refinement strategies applied to each dataset to reduce their AQE. “+ type” refers to refinements related to question types, while “+ domain” refers to refinements based on question domains, following Table 10
Mintaka
Mintaka is a challenging multilingual QA dataset consisting of 20,000 question–answer pairs collected from MTurk contributors and annotated with corresponding Wikidata entities for both questions and answers. Mintaka includes five types of question–answer pairs (entity, boolean, numerical, date, and string) and eight categories (movies, music, sports, books, geography, politics, video games, and history). Among multiple languages, we only use English question-answer pairs.
(1) + type : Among five types, we exclude boolean and numerical questions, following the discussion in §3.1 and §B. (2) + domain : We randomly selected half of the domains (books, movies, music, sports) as the training and validation sets, and assigned the remaining domains to the test set.
ParaRel
ParaRel is originally a dataset designed for masked language modeling, containing factual knowledge expressed through diverse prompt templates and relational types. We utilize the rearranged version by [42]. This version consists of 25,133 prompt-answer pairs across 31 domains. It is further divided into two parts: the first 15 domains are classified as in-domain data, and the remaining 16 domains are classified as out-of-domain.
(1) + domain: In the original setting, in-domain data was used as the test set. In the refined setting, the test set consists of out-of-domain data.
HotpotQA
HotpotQA is a question-answering dataset where each instance consists of a question, label (types including entity, boolean, numerical), and reference documents. We utilize only the question and answer to fit the closed-book scenario. An example of the question is “What government position was held by the woman who portrayed Corliss Archer in the film Kiss and Tell?”, paired with the label “Chief of Protocol”. We use the development dataset as a test set, following [42].
(1) + type: We exclude the “comparison” type, as it consists of yes/no questions or questions that require choosing between two given options.
Explain
We present a benchmark Explain to evaluate a model’s ability to provide a descriptive answer to an open-ended question. Explain is an extended and refined version of an open-ended long-form dataset in the well-known and verified work of FActScore [27]. In FActScore, a small dataset is devised to test the fact-checking pipeline for long-form QA. This dataset is created by appending prompts like “Tell me a bio of <entity>” to person names sourced from Wikipedia. However, its subjects are limited to only person names, and it includes only 500 entries.
To address this, we developed Explain. Explain covers more general categories such as people, history, buildings, culture, etc (the entities from Mintaka), with the dataset size expanded to about 15000 entries. The prompt is "Please give me an explanation about <entity>", which follows the concept of the dataset in FActScore. All entities used as objects in the Explain setting are sourced from entity-type questions in the Mintaka dataset.
(1) + domain: Since the entities in Explain are sourced from Mintaka, we adopt the same domain-splitting strategy as used in Mintaka.
D.2 Data statistics
We present data statistics of our main benchmarks, Mintaka and ParaRel. We also present examples of questions, categories, and statistics on Explain (Table 8, Table 6)
Table 6: #data in each benchmarks
ParaRel Mintaka HaluEval HotpotQA Explain
Train 5575 7583 6000 8000 7583
Valid 5584 1075 2000 2000 1075
Test 13974 2152 2000 7405 2152
Table 7: Examples of questions in Explain
Questions Entity
Please give me an explanation about “A Game of Thrones”. A Game of Thrones
Please give me an explanation about “Simone Biles”. Simone Biles
Please give me an explanation about “Winston Churchill”. Winston Churchill
Please give me an explanation about “Fyodor Dostoevsky”. Fyodor Dostoevsky
Please give me an explanation about “District 12”. District 12
Please give me an explanation about “The Battle of Gettysburg”. The Battle of Gettysburg
Table 8: #Data for the entity domains in Explain
Train Dev Test
Music 914 139 273
History 1059 149 296
Geography 1033 144 306
Politics 1036 143 300
Video games 1057 150 302
Movies 953 138 269
Books 1020 140 283
Sports 909 128 245
Appendix E Experimental detail
E.1 Detail on the three main approaches
In this section, we provide detailed descriptions of the three main approaches evaluated in §5. (1) confidence-based: We adopt a simplified method that takes the top-$n$ softmax probabilities of the first answer token and applies a threshold. $n$ and the threshold $t$ are learnable $\phi$.
For the threshold-based discrimination, we first use the mean value of top-$n$ vocabulary ($s_{j}$ for $j_{th}$ token confidence in top-$n$ candidates) and apply the threshold, as depicted in Equation 5. Here, the learnable parameter $\phi=\{t,n\}$ consists of a threshold ($t$) and the number of vocabulary candidates ($n$). During the training session, every possible pair of $n$ and threshold ($n$ is 1 to 30 in 30 steps, $t$ is 0 to 0.1 in 3000 steps, total 90K $\{t,n\}$ pairs) are measured on the training dataset, and the pair with the highest accuracy is applied to the test session.
$\phi(s)=\begin{cases}1,&\text{if }\frac{1}{n}\sum_{j=1}^{n}s_{j}\geq t\\ 0,&\text{if }\frac{1}{n}\sum_{j=1}^{n}s_{j}<t\par\end{cases}$ (5)
Figure 8: Structure of aggregation of hidden-state and confidence scores.
(2) hidden-state-based: We employ a 3-layer deep neural network (DNN) structure with dimensions $d\rightarrow 100\rightarrow 30\rightarrow 1$, where $d$ is the hidden size. ReLU activation is applied between each layer. The objective function of DNN is binary cross entropy loss $L=-\frac{1}{N}\sum[y\cdot\log(\phi(s))+(1-y)\cdot\log(1-\phi(s))]$. DNN ($\phi$) is trained on the dataset while $\theta$ is frozen. The choice of which layer’s hidden state from $\theta$ to use is determined during the training phase, based on the one that achieves the highest validation accuracy. This approach extends the linear probing [19].
We analyze that DNN emulates the mechanism of the mean threshold approach. The weights of the first layer decide how many candidates to count in, corresponding to the function of $n$ in the threshold-based approach. The second layer decides operations, such as mean or max pooling. DNN structure is a more suitable choice if feature fusion with other data is required.
(3) aggregation: We utilize the feature fusion of confidence and hidden state. That implies utilizing both top-30 confidence value and $h_{th}$ hidden state from $\theta$ as inputs to DNN. This approach concatenates the confidence scores and hidden state into a single vector, which is then passed to a learnable $\phi$ [35]. We use a module with dimensions of $(d+n)\rightarrow 100\rightarrow 30\rightarrow 1$, where $n$ is fixed to 30 (Figure 8).
E.2 Experiment pipeline
First the dataset is divided into $D_{train}$, $D_{valid}$, and $D_{test}$. We fit $\phi$ to $D_{train}$, while $\theta$ is frozen. The next step varies between the two types.
Learning-based
The methods with hidden-state (Probe and Probe + Conf) should employ DNN architecture for $\phi$, which need machine learning. In this case, $\phi$ is trained on the $D_{train}$ with the objective of BCELoss. We train for 20 epochs to get $\phi$. And from the final checkpoint, we choose the index of the hidden layer of $\theta$ with the best accuracy on the $D_{valid}$. Then we use this hidden layer index and $\phi$ to test on the $D_{test}$. We calculate two metrics of accuracy and AUROC. When training, the learning rate is 1e-3, and the optimizer is AdamW.
Threshold-based
The method that solely relies on confidence score (Conf) uses a threshold for calibration. Here, learnable $\phi$ is the number of top candidate confidence scores ($n$) and the threshold ($t$). These two parameters are fitted in $D_{train}$, without evaluation on $D_{valid}$. We select the $\phi$ that achieves the highest accuracy by performing a grid search over $t$ values in the range $[0,1]$ with 1000 uniformly spaced points, and $n$ in the range $[1,30]$ with 3000 uniformly spaced points. And use this $\phi$ to test on the $D_{test}$. AUROC is measured only with $n_{\phi}$, without $t_{\phi}$.
E.3 AQE of larger model
In this section, we provide the AQE results of the larger model (LLaMA-3-70B-Instruct) on the refined dataset, corresponding to Table 9.
Table 9: AQE score of dataset and LLaMA3-70B-Instruct. The version (original, type , domain) with the lowest AQE within each dataset is highlighted in bold.
Mintaka HotpotQA ParaRel Explain
original + type + type + domain original + type original + domain original + domain
$p(k=True)$ 62.17 57.28 57.42 33.81 26.64 51.71 53.01 55.71 54.35
$p(k=False)$ 37.82 42.71 42.57 66.18 73.35 48.28 46.98 44.28 45.64
AQEacc 65.52 62.15 58.96 68.06 71.44 76.68 53.22 61.96 55.31
AQEauc 65.75 64.35 58.77 63.78 54.56 85.98 53.29 67.69 57.71
E.4 Experiment with accuracy
In this section, we present the performance and accuracy-based deltas of various hallucination prediction methods. The results exhibit trends similar to those observed when using AUROC as the evaluation metric.
Table 10: Hallucination prediction performance (accuracy) of instruction-tuned 8B and 70B LLaMA models across multiple datasets.
8B 70B
original + type + type + domain original + type + type + domain
auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$ auroc $\mathcal{A}(\phi(s_{M}))$
Conf 62.75 - 60.71 - 61.54 - 68.39 - 65.53 - 64.65 -
Conf (SCAO) 67.35 - 65.78 - 66.75 - 70.53 - 68.15 - 66.98 -
Probednn 70.55 7.05 68.87 9.06 66.68 7.64 73.32 7.8 69.73 7.58 66.12 7.16
Conf + Probe 71.38 7.88 69.98 10.17 65.12 6.08 74.21 8.69 71.44 9.29 65.99 7.03
Conf + Probe (SCAO) 71.96 8.46 79.41 10.68 68.21 9.17 74.14 8.62 71.35 9.20 67.61 8.65
(a) Mintaka
(b) HotpotQA
8B 70B
original + type original + type
acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$
Conf 71.54 - 75.93 - 69.75 - 73.08 -
Conf (SCAO) 73.23 - 76.93 - 70.97 - 75.97 -
Probednn 75.49 6.94 76.31 0.28 71.24 3.18 71.91 0.47
Conf + Probe 76.00 7.45 75.46 -0.56 72.50 4.44 74.02 2.58
Conf + Probe (SCAO) 77.69 9.14 76.95 0.92 72.51 4.45 74.90 3.46
(c) ParaRel
8B 70B
original + domain original + domain
acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$
Conf 67.58 - 62.63 - 65.75 - 57.29 -
Conf (SCAO) 66.67 - 67.88 - 66.83 - 69.19 -
Probednn 80.52 7.26 66.82 11.73 82.29 5.61 65.24 12.02
Conf + Probe 80.64 7.38 68.65 13.56 82.18 5.50 64.46 11.24
Conf + Probe (SCAO) 80.92 7.66 69.24 14.15 81.84 5.16 66.36 13.14
(d) Explain
8B 70B
original + domain original + domain
acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$ acc $\mathcal{A}(\phi(s_{M}))$
Conf 57.66 - 60.29 - 55.71 - 53.31 -
Conf (SCAO) 59.71 - 59.57 - 60.99 - 60.78 -
Probednn 78.92 13.27 63.31 2.1 76.28 14.32 62.65 7.34
Conf + Probe 79.12 13.47 64.78 3.57 76.00 14.04 63.10 7.79
Conf + Probe (SCAO) 80.91 15.26 63.66 2.45 78.75 16.79 64.87 9.56
Appendix F Instruction prompts
In this section, we compile the instructional prompts employed in our study. Terms marked with $\underline{\texttt{underline}}$ indicate placeholders that need to be filled with the corresponding content.
A. Normal instruction template
[Question]:{question} [Answer]:
B. SCAO instruction template
[Question]: {question} You must answer in only one word. [Answer]:
C. G-eval instruction template
[Instruction] The text enclosed within triple backticks (“‘) is an agent’s answer to a given question ([Question]). Your task is to evaluate whether the agent’s response is factually correct. . 1) Analyze and explain whether the answer contains any factual inaccuracies. 2) Then, classify the answer as either "True" (only factually correct contents) or "False" (containing any factually incorrect content). . “‘ [Question]:{question} [Answer]:{answer} ”’